# pnasresearcharticle

### v1

@@ Abstract
Extreme moist heatwaves pose a serious threat to our society and human health. To manage heat-related risks, it is crucial to improve our understanding of what sets the maximum moist heat. This question has been extensively studied for moist heat in the tropics under the assumption of moist neutrality, but it remains less explored over midlatitude continents, where this assumption is not applicable. The frequent concurrent occurrence of extreme moist heat and intense convection over midlatitude land supports the prevailing hypothesis that deep convection terminates heatwaves and sets their peak intensity. Here a theory is eveloped based on the onset of convection to provide a theoretical prediction for both the maximum near-surface moist heat and potential convection over midlatitude land. Our framework demonstrates that the intensities of moist heat and convection are constrained by a preexisting energy barrier in the lower free troposphere, which suppresses convection and enables the buildup of moist heat and convective instability. This barrier, often marked by a temperature or energy inversion, varies only slightly over time, suggesting the potential to predict maximum intensities of moist heat and convection from predefined lower free tropospheric properties. This framework aligns with the stored-energy nature of deep convection over midlatitude continents, offering insights into understanding the evolution of extreme heat and convection.
keywords:
atmospheric stability moist heat deep convection convective inhibition energy inversion
@@
Moist heatwaves, which combine the effects of high temperature and humidity, pose significant risks to human health (1). While previous research has primarily focused on the dynamics and changes of moist heat in tropical and subtropical regions (2, 3, 4), the drivers and characteristics of extreme moist heat over midlatitude continents remain poorly understood. As climate change is driving severe moist heat risks poleward into the midlatitudes (5), advancing our theoretical understanding of the limits of maximum moist heat in these regions has become increasingly urgent. Our goal in this study is to determine what sets the maximum moist heat over midlatitude land.
@@
Near-surface wet bulb temperature (WBTs) is a key measure of moist heat stress, as it combines the influence of temperature and humidity (6, 3, 5, 7). By definition, WBT represents the temperature an air parcel would attain if cooled adiabatically to saturation at constant pressure by evaporating water into it. During this adiabatic process, the air parcel conserves its moist static energy (MSE) or the equivalent potential temperature, and thus WBTs can be directly derived assuming the conservation of near-surface MSE (MSEs). Unlike the non-linear WBT formula, MSE is a linear function of temperature, specific humidity, and geopotential height, making it simple to use. Accordingly, this study employs MSEs, defined as the sum of sensible heat, latent heat, and geopotential energy at 2 meters above the ground, to represent moist heat, to measure moist heat. The MSEs reasonably represents moist heat, as the annual maximum MSEs (Fig. 1) recovers well the spatial distribution of moist heat indicated by the annual maximum WBTs (Fig. 1). There is a small variation in WBTs given a value of MSEs, due to the dependence on the elevation variation (0β1000 m; Fig. S1). In this work, we calculated MSEs and other variables analyzed using the 3-hourly ERA5 reanalysis data during 1980β2022 for midlatitude land grids between 35βN and 75βN with elevation lower than 1000 m. Our analysis focuses on the annual maximum MSEs to explore the potential constraint on the maximum MSEs.
@@
Regarding the maximum MSEs, previous studies have commonly assumed a strong coupling between the near-surface atmosphere and the free troposphere, such that MSEs is constrained by the free troposheric (typically 500-hPa) saturated moist static energy (MSE) through deep convection. Under the assumption of a moist-neutral atmospheric column, this framework has been widely used to study extreme heat in the tropics (3, 2). With the same neutrality assumption, recent studies (8) have applied this framework to investigate extreme dry heat over midlatitude continents, suggesting that MSEMSE. We first test this existing theory by comparing MSEs with MSE during the time of annual maximum moist heat for all land grids in the midlatitudes. Our analysis reveals that this theory does not hold for extreme moist heat, as MSEs is consistently and significantly greater than MSE in nearly all cases (Fig. 1). Furthermore, the distribution of MSE is largely zonally uniform (Fig. S1), which does not capture the strong zonal variation observed in the MSEs pattern(Fig. 1).
@@
Here, we build on prior studies to propose a new theoretical framework that accounts for the onset of convection, which enables surface heat and convective instability to evolve beyond the neutral condition, and ultimately provides a more tight constraint on near-surface moist heat over midlatitude land.
Concurrent occurrence of moist heat and convection extremes
@@
We first demonstrate that deep convection occurs concurrently with the annual maximum moist heat (Fig. 2), supporting the hypothesis that convective instability acts to terminate heatwaves and constrain maximum near-surface heat. The potential intensity of convection is measured by convective available potential energy (CAPE; detailed in ), with values of 1000 J kg-1 or higher typically deemed sufficient to support strong to severe convection, according to National Weather Service guidelines (9). We define the CAPE at the time of the annual maximum MSEs as critical CAPE (denoted CAPEc), which measures the potential convection that terminates the annual maximum moist heat. The CAPEc over most midlatitude land regions exceeds 1000 J kg-1, with values ranging from 3000 to 4000 J kg-1 across western Europe extending to Northeast China, and reaching 6000 J kg-1 or higher over eastern China and central North America (Fig. 2). The high values of CAPEc represent substantial convective instability in the atmosphere during extreme moist heat, indicating a significant departure from the neutrality assumption. Meanwhile, critical convective inhibition (denoted CINc), a measure of the energy barrier in the lower free troposphere that suppresses the release of CAPE (detailed in ), is sufficiently low to allow the initiation of free convection (Fig. S2).
@@
In fact, the deep convection that terminates the maximum moist heat is likely the most intense convection of the year across many midlatitude land areas, as CAPEc (Fig. 2) closely aligns with the annual maximum CAPE (Fig. S2). This alignment is particularly evident in Europe and central North America, where the difference between CAPEc and the annual maximum CAPE is less than 250 J kg-1 (Fig. S2). In these regions, over 50% of the years between 1980 and 2022 feature simultaneous annual maxima of CAPE and MSEs (Fig. S2), emphasizing the frequent co-occurrence of extreme moist heat and intense convection in the midlatitudes.
@@
We further provide a scaling of CAPE in terms of MSEs to demonstrate why both moist heat and convection maximize at the same time. Based on definitions of parcel buoyancy and simple assumptions about the geometry of CAPE, CAPE can be approximated by the difference between MSEs and 500-hPa saturated moist static energy (MSE) scaled by a factor of 0.22 (detailed in ):
@@
The scaling CAPE builds a simple linear relation between CAPE and basic environmental parameters of the near-surface and free tropospheric atmospheres, without relying on the profiles of a hypothetically lifted air parcel. The scaling CAPEc, estimated using MSEs and MSE at the time of yearly maximum MSEs, effectively captures the spatial distribution of CAPEc (Fig. 2 and ), with values closely aligned along the one-to-one line (Fig. 2). We assume that MSE remains relatively constant over short time periods, which is consistent with previous studies suggesting that the free troposphere experiences minimal variation during the heatwave evolution (8) or convection build-up periods (10, 11). Consequently, based on Eq.1, CAPE are theoretically expected to evolve linearly with MSEs and reach their peaks simultaneously. The frequent concurrence of extreme near-surface moist heat and convective instability suggests a shared governing process that drives the accumulation of both phenomena over midlatitude land. This implies that the factor determining the maximum MSEs simultaneously sets the maximum CAPE. We proceed to provide a theoretical prediction for both moist heat and convection extremes.
Theoretical framework
@@
We stick to the hypothesis that the onset of deep convection sets the maximum potential intensity of both near-surface moist heat and convection that can be achieved, but we incorporate a more detailed consideration of convection onset to propose a new theoretical framework. This framework allows surface heat and convective instability to evolve beyond the neutral condition, providing a tight constraint on both near-surface moist heat and convection throughout this process.
@@
The foundation of our framework is the stored-energy nature of severe convection over midlatitude continents (12, 13), primarily attributed to a preexisting stable lower free troposphere (14, 10, 11). For instance, as shown in a sample sounding from a week prior to the 2014 yearly maximum surface moist heat over the central United States (temperature profiles in Fig. 3), the low-level atmospheric stability typically creates a strong layer of negative buoyancy (i.e., < 0 and CIN > 0) that inhibits convective initiation, even when the free troposphere may already be unstable (i.e., > 0 and CAPE > 0). Therefore, assuming sufficient heat sources, the near-surface atmosphere can continue to warm or moisten until the negative buoyancy layer is eroded, enabling deep free convection to initiate, release CAPE, and cool the surface. In the shown case (temperature profiles in Fig. 3), near-surface heat increased over subsequent days, accompanied by a rise in CAPE, while the free troposphere remained relatively constant, until the the time of maximum surface heat when the convective inhibition was nearly eliminated. Eventually, at the critical point when both the near-surface moist heat (MSEs) and the potential intensity of convection (CAPE) reach their maximum values, the local atmospheric environment is characterized by CINc = 0 and CAPEc 0, consistent with the results presented above (Fig. S2 and 2).
@@
Therefore, the regime of near-surface moist heat accumulation and convection buildup is constrained by CIN 0. Since CIN essentially reflects the presence of negative buoyancy layers within the lower atmosphere, whether and when CIN approaches 0 is governed by the evolution of the most negative buoyancy () throughout the low levels. Therefore, CIN 0 is effectively equivalent to 0. The latter is more general and practical, as it also captures cases where CIN is 0 by definition due to the absence of a defined level of free convection (LFC) and equilibrium level (EL) with the entire atmospheric column characterized by negative buoyancy for the air parcel.
@@
As parcel buoyancy is proportional to the temperature difference between the air parcel and its environment, we further utilize the conservation of the parcelβs MSE to derive an alternative formulation of temperature perturbations entirely as a function of environmental static energies (Eq.8, detailed in ). This also provides an MSE perspective to visualize the profiles of temperature perturbations and buoyancy (Eq.9, detailed in ) by comparing the parcelβs initial MSE (i.e., ) with the environmental static energy profile ( if or if , where represents the lifted condensation level). As is usually above the LCL (11), and thus based on Eq.9, 0 is equivalent to
@@
As MSEs does not change with height, = MSEMSE, where MSE is the maximum MSEβ over the layers above or at LCL. Therefore, the maximum is limited by MSE, such that
@@
From the equation of scaling CAPE (Eq.1), the MSE and MSE together constrain the CAPEc by
@@
The right-hand side of Eqs.3 and 4 defines the maximum potential intensities of near-surface moist heat (MSEs) and convection (CAPE), respectively. These maximum potential intensities are achievable under conditions where surface heat sources are sufficient and external lifting forces are absent, requiring to approach zero to enable free convection, which is commonly observed in midlatitude continents (10, 11). Otherwise, the actual intensities are likely to be lower than their maximum potential values.
@@
We first present the results of the sounding evolutions for the same case discussed above, analyzed within the developed MSE framework. As a validation of Eq.9, the MSE profiles (Fig. 3 and ) recover well the positive and negative buoyancy layers and their temporal evolution. Compared with temperature profiles, the MSE profiles clearly characterize the most negative buoyancy layer prior to the yearly maximum moist heat (MSE profiles in Fig. 3), identified by an explicit MSEβ inversion above the LCL. Compared to temperature profiles, the MSE profiles more clearly delineate the most negative buoyancy layer prior to the annual maximum moist heat, identified by a distinct MSEβ inversion above the LCL (Fig. 3). The preexisting MSE resides at the top of the MSEβ inversion near 800 hPa and eventually aligns with the LCL (MSE profiles in Fig. 3 and ). In this case, the final MSE effectively sets the maximum MSEs achieved at the time of the annual maximum moist heat, when (and hence CINc) approaches 0 and CAPEc is maximized. Relative to the significant changes in MSEs during this period, the preexisting MSE intensities only slightly over time, which suggests the possibility of predicting the maximum intensity of MSEs (and consequently CAPEc) based on the initial MSE, although adjustments are needed over time as MSE evolves.
@@
To evaluate the generality of the results from the case study, we extend the analysis by presenting composite time series of key proxies derived from a broader set of cases over midlatitude land. Specifically, we focus on instances of concurrent annual maxima in MSEs and CAPE for each year during 1980β2022. To best align with our theoretical framework - where MSEs and CAPE are maximized as preexisting high CIN is eroded to zero - we further filter cases based on the following criteria: CINc = 0, CAPEc > 500 J kg-1, and a maximum CIN of at least 150 J kg-1 within the seven days preceding the critical time. These selection criteria yield 535 cases, for which we plot composite 14-day time series centered on the time of maximum MSEs (Fig. 3β). Following the maximum potential for moist heat defined by Eq.3, MSEs increases over time and ultimately converges toward MSE at the peak time (Fig. 3). Notably, MSE fails to predict the maximum MSEs as discussed above, as MSEs exceeds MSE as early as seven days before the peak, with the difference rapidly amplifying over time (Fig. 3). This growing discrepancy between MSEs and MSE reflects the progressive accumulation of CAPE leading up to the peak, as recovered by the scaling CAPE following Eq.1 and constrained by the maximum potential for convection defined in Eq.4 (Fig. 3). Fundamentally, the composite temporal evolutions of MSE and CAPE are linked to the removal of CIN or (Fig. 3), supporting our basic hypothesis formulated in Eq.2. Similarly, the composite results reveal relatively steady time series for MSE and MSE (Fig. 3), as well as for the potential CAPE (Fig. 3), highlighting the potential to predict the maximum intensities of MSEs and CAPE based on preexisting free tropospheric properties.
@@
Lastly, to comprehensively validate the potential intensity framework (i.e., Eqs.3 and 4), we compare the observed annual maximum moist heat (MSEs) and the associated critical convection (CAPEc) with their potential maxima derived for all midlatitude land grids during 1980β2022 using the ERA5 reanalysis data (Fig. 4). Note that our goal here is not to evaluate the predictive capability of our theory based on prior atmospheric properties but rather to provide a precise, apple-to-apple assessment at the peak time. Accordingly, MSE and MSE are calculated at the time of the annual maximum MSEs. Overall, the potential MSEs (Fig. 4) and the potential CAPEc (Fig. 4) well capture the spatial patterns of the observed maximum MSEs (Fig. 1) and CAPEc (Fig. 2) over midlatitude land, with observed maxima only slightly smaller than their potential counterparts. The pattern correlation coefficient is 0.94 between observed and potential MSEs, and 0.81 between observed and potential CAPEc. The joint histograms of all cases further confirm that the potential intensities serve as a tight upper limit for the observed maximum MSEs and CAPEc (Fig. 4 and ). For the maximum MSEs, the potential MSEs is achievable for many cases, with others closely aligned with the one-to-one line but slightly below (Fig. 4). Similarly, the potential CAPEc acts as a tight constraint for the observed CAPEc as well (Fig. 4), though the variation in observed CAPEc for a given potential CAPEc appears larger, partly due to the variability introduced by the scaling of CAPEc (Fig. 2)
Conclusions and discussions
This study proposed a theoretical prediction for maximum moist heat (MSEs) and convection (CAPE) over midlatitude land, both tightly constrained by a preexisting energy (MSEβ) inversion within the lower free troposphere. The maximum MSEβ at the energy inversion layer (i.e., MSE) acts as a substantial energy barrier that suppresses the free convection of near-surface air parcels, thereby allowing the warming or moistening of near-surface air and the accumulation of convective instability. This process continues until MSEs reaches MSE, such that the near-surface air parcels are sufficiently energetic to overcome the energy barrier and initiate deep free convection. Recent studies have explored a similar topic but for dry heat extremes under the assumption of moist neutrality (8), suggesting that the 500-hPa MSE serves as an upper bound for MSEs. However, this assumption fails for extreme moist heat as MSEs generally exceeds MSE during the maximum moist heat period in midlatitudes. Notably, our scaling CAPE framework demonstrates that the surplus of MSEs relative to MSE effectively defines the potential convection intensity of the environment. This study extended prior work by carefully addressing the onset of convection, which enables the atmosphere to evolve beyond the neutral condition, consistent with the stored-energy nature of severe convection over midlatitude continents. Future work may build upon our work to incorporate other processes (e.g., wind shear, entrainments, and external forcings) to more precisely determine the onset of convection and, consequently, the termination of heatwaves. For example, recent research has highlighted the critical role of dry air entrainment in suppressing deep convection and intensifying moist heat in tropical and subtropical regions (4), though the entrainment influence might be less important in midlatitude continental convection due to the generally strong vertical wind shear (15).

### v3

# Atmospheric stability sets maximum moist heat and convection in the midlatitudes

@@ Abstract
Extreme moist heatwaves pose a serious threat to society and human health. To manage heat-related risks, it is crucial to improve our understanding of what limits the maximum near-surface moist heat. This question has been extensively studied for moist heat in the tropics under the assumption of moist neutrality, but it remains less explored over midlatitude continents, where this assumption is not applicable. The frequent concurrent occurrence of extreme moist heat and intense convection over midlatitude land reinforces the prevailing hypothesis that deep convection terminates heatwaves and sets their peak intensity. Here, a theory is developed based on the onset of convection to provide a theoretical prediction for both the maximum near-surface moist heat and potential convection over midlatitude land. Our framework demonstrates that the intensities of moist heat and potential convection are constrained by a preexisting energy barrier in the lower free troposphere, which suppresses convection and enables the buildup of moist heat and convective instability. This barrier, often marked by a temperature or energy inversion, varies only slightly during moist heatwave evolution. Our results suggest the potential to predict maximum moist heat and convection from predefined lower free tropospheric properties and offer insights into the evolution of extreme heat and convection across climate states.
This manuscript was compiled on January 22, 2025
@@
Moist heatwaves, which combine the effects of high temperature and humidity, pose significant risks to public health and societal outcomes (1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12). While previous research has primarily focused on the dynamics and changes of moist heat in tropical and subtropical regions (13, 14, 15, 16, 17, 18, 19), the drivers and characteristics of extreme moist heat over midlatitude continents remain poorly understood. As climate change is driving severe moist heat risks poleward into the midlatitudes (20, 21, 22, 23, 24, 25, 26, 27), advancing our theoretical understanding of the limits of maximum moist heat in these regions has become increasingly urgent. Our goal in this study is to determine what sets the maximum moist heat over midlatitude land.
@@
Near-surface wet bulb temperature (WBTs) is a key measure of moist heat stress, as it combines the influence of temperature and humidity (28, 3, 29, 4, 23, 16, 26, 30). By definition, WBT represents the temperature that an air parcel would attain if cooled adiabatically to saturation at constant pressure by evaporating water into it (28, 31, 32). During this adiabatic process, the air parcel conserves its moist static energy (MSE) or the equivalent potential temperature, and thus WBTs can be directly derived assuming the conservation of near-surface MSE (MSEs) (33, 16, 15, 30, 34). Unlike the non-linear WBT formula, MSE is a linear function of temperature, specific humidity, and geopotential height, making it simple to use (15, 35). Accordingly, this study employs MSEs, defined as the sum of sensible heat, latent heat, and geopotential energy at 2 meters above the ground, to measure near-surface moist heat. Overall, MSEs reasonably represents moist heat and, similar to the annual maximum WBTs (Fig. 1), the annual maximum MSEs recovers well the spatial distribution of extreme moist heat over midlatitude land (Fig. 1) and aligns closely with WBTs (Fig. S1; detailed in ). In this work, we use the 3-hourly ERA5 reanalysis data (36, 37) during 1980–2022 for land grids between 35∘N and 75∘N with elevation lower than 1000 m (detailed in ). Our analysis focuses on the annual maximum MSEs to explore the potential constraint on the maximum moist heat over midlatitude land.
@@
Previous studies have commonly assumed a strong coupling between the near-surface atmosphere and the free troposphere during extreme heat through deep moist convection, such that MSEs is constrained by free troposheric saturated moist static energy (MSE∗) given the state of convective quasi-equilibrium (38, 39, 40, 3, 41, 42, 43, 44, 16, 14, 4). Under the assumption of a moist-neutral atmospheric column, this framework has been widely used to study extreme heat in the tropics (14, 16, 18) or the global mean state (4). With the neutrality assumption, a recent study (45) has applied this framework to investigate extreme dry heat over midlatitude continents, suggesting that MSEs is limited by 500-hPa MSE∗ (i.e., MSEMSE), thereby setting an upper bound for maximum surface air temperature. We first test the validity of this theory by comparing MSEs with MSE during the annual maximum moist heat over midlatitude land. Our analysis reveals that the theory does not hold for extreme moist heat, as MSEs is consistently and significantly greater than MSE in nearly all cases (more than 96% of the midlatitude land grid points; Fig. 1–). Furthermore, the distribution of MSE is largely zonally uniform (Fig. 1), which does not capture the strong zonal variation observed in the MSEs pattern (Fig. 1).
@@
These results indicate a violation of the assumption of moist neutrality because, unlike quasi-equilibrium convection in the tropics, severe convection over midlatitude continents involves the strong storage of potential energy over a short time period (46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56), which potentially interacts with surface heat in distinct ways. Here, we propose a new theoretical framework that accounts for the onset of convection, which enables surface heat and convective instability to evolve beyond the neutral condition, and ultimately provides a tight constraint on both near-surface moist heat and potential convection over midlatitude land.
Concurrent moist heat and convection extremes
@@
We first demonstrate that deep convection occurs concurrently with the annual maximum moist heat (Fig. 2), supporting the hypothesis that convective instability acts to terminate heatwaves and constrain maximum near-surface heat (4, 45, 57, 58, 17). The potential intensity of convection is measured by convective available potential energy (CAPE; detailed in ), with values of 1000 J kg-1 or higher typically deemed sufficient to support strong to severe convection, according to the National Weather Service guidelines (59). We define CAPE at the time of the annual maximum MSEs as critical CAPE (denoted CAPEc), which measures the potential convection that terminates the annual maximum moist heat. The CAPEc over most midlatitude land regions exceeds 1000 J kg-1, with values ranging from 3000 to 4000 J kg-1 across western Europe extending to Northeast China, and reaching 6000 J kg-1 or higher over eastern China and central North America (Fig. 2). The high values of CAPEc represent substantial convective instability in the atmosphere during extreme moist heat, indicating a significant departure from the neutrality assumption. Meanwhile, critical convective inhibition (denoted CINc; detailed in ), a measure of the energy barrier in the lower free troposphere that suppresses the release of CAPE, is sufficiently low (generally less than 25 J kg-1) to allow the initiation of deep free convection in the midlatitudes (Fig. S2).
@@
In fact, the deep convection that terminates the maximum moist heat is likely the most intense convection of the year across many midlatitude land areas, as CAPEc (Fig. 2) aligns closely with the annual maximum CAPE (Fig. S2). This alignment is particularly evident in Europe and central North America, where the difference between CAPEc and the annual maximum CAPE is less than 250 J kg-1 (Fig. S2). In these regions, over 50% of the years during 1980–2022 feature simultaneous annual maxima of CAPE and MSEs (Fig. S2), emphasizing the frequent co-occurrence of extreme moist heat and intense convection in the midlatitudes.
Theoretical basis
CAPE scaling
@@
We start with providing a scaling of CAPE in terms of MSEs to demonstrate why both moist heat and potential convection maximize at the same time. Based on the definition of parcel buoyancy and simple assumptions about the geometry of CAPE, CAPE can be approximated by the difference between MSEs and MSE scaled by a factor of 0.22 (detailed in ):
@@
The scaling CAPE (right hand side of Eq.1) builds a simple linear relation between CAPE and basic environmental parameters of the near-surface and free tropospheric atmospheres, without relying on the profiles of a hypothetically lifted air parcel. The scaling CAPEc, estimated using MSEs and MSE at the time of annual maximum MSEs, effectively captures the spatial distribution of CAPEc (pattern correlation coefficient = 0.98; Fig. 2 and ), with values closely aligned along the one-to-one line (Fig. 2). We assume that MSE remains relatively constant over short time periods, which is consistent with previous studies suggesting that the free troposphere experiences minimal variations during the heatwave evolution (60, 45) or convection build-up periods (55, 56). Consequently, based on Eq.1, CAPE and MSEs are linearly proportional and expected to reach their peak intensities simultaneously. The frequent concurrence of extreme near-surface moist heat and convective instability suggests a shared governing process that drives the accumulation of both phenomena over midlatitude land, thereby the factor determining the maximum MSEs simultaneously sets the maximum CAPE. We proceed to provide a theoretical prediction for both maximum moist heat and maximum potential convection.
A theory for convection onset
@@
We stick to the hypothesis that the onset of deep convection sets the maximum potential intensity of both near-surface moist heat and convection, but we incorporate a more detailed consideration of convection onset to propose a new theoretical framework. This framework allows surface heat and convective instability to evolve beyond the neutral condition, providing a tight constraint on both near-surface moist heat and convection throughout the process.
@@
The foundation of our framework is grounded in the stored-energy nature of severe convection over midlatitude continents (46, 47, 48, 49, 50, 52, 53, 54), primarily attributed to a preexisting stable lower free troposphere (61, 62, 63, 15, 64, 65, 55, 56, 66). For instance, as shown in a sample sounding from a week prior to the annual maximum surface moist heat in 2004 over the central United States (temperature profiles in Fig. 3), the low-level atmospheric stability typically creates a strong layer of negative buoyancy (i.e., < 0 and CIN > 0) that inhibits convective initiation, even when the free troposphere may already be unstable (i.e., > 0 and CAPE > 0). Therefore, assuming sufficient heat sources, the near-surface atmosphere can continue to warm or moisten until the negative buoyancy layer is eroded, enabling the initiation of deep free convection to release CAPE and terminate the surface heat through precipitation followed by evaporative cooling. In the case study presented, near-surface heat progressively increased over the following days, accompanied by a rise in CAPE. Meanwhile, the free troposphere remained relatively constant until the time of maximum surface moist heat, when convective inhibition was nearly eliminated (temperature profiles in Fig. 3). At the critical point, when both the MSEs and CAPE reach their maximum values, the local atmospheric environment is characterized by CIN 0 and CAPEc 0 (temperature profiles in Fig. 3), consistent with the results presented above (Fig. S2 and 2).
@@
Therefore, the regime of near-surface moist heat accumulation and convection buildup is constrained by CIN 0. Since CIN essentially reflects the presence of negative buoyancy layers within the lower atmosphere, whether and when CIN approaches 0 is governed by the evolution of the most negative buoyancy () throughout the low levels (55, 56). Therefore, CIN 0 is effectively equivalent to 0. The latter is more general and practical, as it also captures cases where CIN is 0 by definition due to the absence of a defined level of free convection (LFC) and equilibrium level (EL), with the entire atmospheric column characterized by negative buoyancy for the near-surface lifted air parcel.
@@
As parcel buoyancy is proportional to the (virtual) temperature difference between the lifted air parcel and its environment, we further utilize the conservation of the parcel’s MSE to derive an alternative formulation of temperature perturbations entirely as a function of environmental static energies (Eq.8, detailed in ). This also provides an MSE perspective to visualize the profiles of temperature perturbations and buoyancy by comparing the parcel’s initial MSE (i.e., ) with the environmental moist static energy profile, which is for or for , where represents the lifted condensation level (Eq.9, detailed in ). As is usually above the (55, 56), based on Eq.9, 0 is equivalent to
@@
As MSEs does not change with height, = MSEMSE, where MSE is the maximum MSE∗ over the lower free tropospheric layers above or at LCL (detailed in ). The MSE quantifies the strongest energy barrier within the lower free troposphere that a near-surface air parcel must overcome to initiate convection in the absence of external forcing. Therefore, the maximum is limited by MSE, such that
@@
In addition, from the framework of scaling CAPE (Eq.1), the MSE and MSE together constrain CAPEc by
@@
The right-hand side of Eqs.3 and 4 defines the maximum potential intensities of near-surface moist heat (MSEs) and convection (CAPE), respectively. These maximum potential intensities are achievable under conditions where surface heat sources are sufficient and external lifting forces are absent, requiring to approach zero to enable free convection, which is commonly observed in midlatitude continents (62, 67, 68, 55, 56). Otherwise, the actual intensities are likely to be lower than their maximum potential values.
@@
We first present the results of the sounding evolutions for the case discussed above, analyzed within the developed MSE framework. As a validation of Eq.9, the MSE profiles (Fig. 3 and ) recover well the positive and negative buoyancy layers and their temporal evolutions. Compared to temperature profiles, the MSE profiles more clearly delineate the most negative buoyancy layer prior to the annual maximum moist heat, identified by a distinct MSE∗ inversion above the LCL (Fig. 3). The preexisting MSE resides at the peak of the MSE∗ inversion near 800 hPa (MSE profiles in Fig. 3), and eventually aligns with the LCL (MSE profiles in Fig. 3). In this case, the final MSE effectively sets the maximum MSEs achieved at the time of the annual maximum moist heat, when (and hence CINc) approaches 0 and CAPEc is maximized. Relative to the significant changes in MSEs during this period, the preexisting MSE intensities only slightly over time, which suggests the possibility of predicting the maximum potential intensity of MSEs (and consequently CAPEc) based on the initial MSE, although adjustments are needed over time as MSE evolves.
@@
To evaluate the generality of the above results, we extend the analysis by presenting composite time series of key proxies derived from a subset of cases over midlatitude land. Specifically, we focus on instances of concurrent annual maxima in MSEs and CAPE for each year during 1980–2022 (44.5% of all cases; Fig. S2). To best visualize the features aligning with our theoretical framework - where MSEs and CAPE are maximized as preexisting CIN is eroded to zero - we further select cases based on the following criteria: CAPEc > 500 J kg-1, CIN 0, and a maximum CIN of at least 50 J kg-1 within seven days preceding the critical time (detailed in ). These selection criteria yield 2502 cases per year (27% of concurrent cases), for which we plot composite 14-day time series centered on the time of maximum MSEs (Fig. 3–). Following the maximum potential for moist heat defined by Eq.3, MSEs increases over time and ultimately converges toward MSE at the peak time (Fig. 3). The height of MSE is in general between 750–800 hPa, and decreases gradually with time to approach (Fig. S3). Notably, MSE fails to predict the maximum MSEs as discussed above, as MSEs exceeds MSE as early as seven days before the peak, with the difference rapidly amplifying over time (Fig. 3). This growing discrepancy between MSEs and MSE reflects the progressive accumulation of CAPE leading up to the peak, as recovered by the scaling CAPE following Eq.1 and constrained by the maximum potential for convection defined in Eq.4 (Fig. 3). Fundamentally, the composite temporal evolutions of MSE and CAPE are linked to the removal of CIN or (Fig. 3), supporting our basic hypothesis formulated in Eq.2. These results are broadly consistent when we perform composite analyses based on all cases of concurrent moist heat and convection over midlatitude land (detailed in ; Fig. S4), though the maximum potential intensities of MSEs and CAPEc are not necessarily achieved (Fig. S4–). This suggests the potential influence of external forcing mechanisms of lift (69, 70, 71), which may facilitate the initiation of convection in cases with non-zero CINc (Fig. S4). Similarly, the composite results reveal relatively steady time series for MSE and MSE (Fig. 3), as well as for the potential CAPE (Fig. 3), highlighting the potential to predict the maximum intensities of MSEs and CAPE based on preexisting free tropospheric properties.
@@
Lastly, to comprehensively validate the maximum potential intensity framework (i.e., Eqs.3 and 4), we compare the observed annual maximum moist heat (MSEs) and the associated critical convection (CAPEc) with their potential maxima derived for all cases over midlatitude land during 1980–2022 from the ERA5 reanalysis data (Fig. 4). Here, MSE and MSE are calculated at the time of annual maximum MSEs. Overall, the maximum potential MSEs (Fig. 4) and the maximum potential CAPEc (Fig. 4) capture the spatial patterns of the observed maximum MSEs (Fig. 1) and CAPEc (Fig. 2) over midlatitude land, with observed maxima slightly smaller than their potential counterparts. The pattern correlation coefficient is 0.94 between observed and potential MSEs maxima, and 0.82 between observed and potential CAPEc. The joint histograms of all cases further confirm that the maximum potential intensities serve as a tight upper limit for the observed maximum MSEs and CAPEc (Fig. 4 and ). For the maximum MSEs, the potential MSEs is achievable for many cases, with others closely aligned with the one-to-one line but slightly below (Fig. 4). Translating (potential) MSEs into (potential) WBTs (as detailed in Materials and Methods) yields consistent results (Fig. 4), highlighting the applicability of our theory for directly predicting maximum WBTs. Similarly, the potential CAPEc serves as a strong constraint on the observed CAPEc (Fig. 4), although the actual CAPEc exhibits greater variability for a given potential CAPEc, partly due to the variations in MSEs.
Conclusion and discussion
This study proposed a theoretical prediction for maximum potential moist heat (MSEs) and convection (CAPE) over midlatitude land, both tightly constrained by a preexisting energy (MSE∗) inversion within the lower free troposphere. The maximum MSE∗ at the energy inversion layer (i.e., MSE) acts as a substantial energy barrier that suppresses the free convection of near-surface air parcels, thereby allowing the warming or moistening of near-surface air and the accumulation of convective instability. This process continues until MSEs reaches MSE, such that the near-surface air parcels are sufficiently energetic to overcome the energy barrier, initiate deep free convection, and terminate surface heating. This theory does not address the mechanisms by which moist heat or convective instability increases, which may be driven by various physical processes (e.g., warm advection, adiabatic warming in subsiding air, or diabatic heating from surface fluxes and radiation) associated with large-scale dynamics such as storm tracks or atmospheric blockings (72, 73, 74, 56, 75, 76, 30, 77, 78, 79, 60, 80). Instead, it focuses on quantifying the upper limit of these increases before free convection is triggered, assuming sufficient heating sources.
A recent study examined a similar topic over midlatitude continents, but focusing on extreme dry heat under the assumption of moist neutrality (45), and suggested that 500-hPa MSE acts as an upper bound for MSEs. However, we have shown that this assumption fails for extreme moist heat over midlatitude land. In fact, the neutrality assumption also breaks down for extreme dry heat in the midlatitudes, as MSEs exceeds MSE in most cases of annual maximum surface air temperature (Fig. S5). Our CAPE scaling framework (Eq.1) reveals that MSE generally approximates the minimum energy of the free troposphere (detailed in ), and the surplus of MSEs over MSE effectively defines the accumulated convective instability beyond moist neutrality. When applied to extreme dry heat conditions, our theory also predicts well the maximum potential intensities of MSEs and CAPEc during extreme dry heat periods (Fig. S5–), thereby generalizing our theory and highlighting the critical role of preexisting low-level energy inversions in setting near-surface maximum heat.
Our study advances previous work by carefully addressing the onset of convection, demonstrating how the atmosphere evolves beyond the neutral condition – where parcel buoyancy and CAPE are zero by definition – in alignment with the stored-energy nature of severe convection in the midlatitudes (46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56). Future work could incorporate other processes (e.g., wind shear (81, 82, 83), entrainments (42, 84, 17), and external forcings (85, 86, 83, 87, 88, 89, 71, 69, 70)) to more precisely determine the onset of convection and, consequently, the termination of heatwaves. For example, a recent study highlighted the critical role of dry air entrainment in suppressing deep convection and intensifying moist heat in tropical and subtropical regions (17), though the entrainment influence may vary strongly with the severity of convection in midlatitude due to the dependence on vertical wind shear (90, 91).